OpenStudy (anonymous):

Simlify the expression: 80^(1/4)/5^(-1/4) Simlify the expression: 80^(1/4)/5^(-1/4) @Mathematics

jhonyy9 (jhonyy9):

80(1/4)/1/5(1/4) = 80(1/4) *5(1/4) = 400(1/4)